OpenStudy (anonymous):

1. Determine whether the sequence converges or diverges. If it converges, give the limit. 60, 10, 5/8 , 5/18 , ... Converges; 72 Converges; 0 Diverges Converges; -15540 2.) find an explicit rule for the nth term of the sequence. 1, 3, 9, 27, ... an = 1 • 3n + 1 an = 3 • 1n - 1 an = 1 • 3n - 1 an = 1 • 3n 3. Find an explicit rule for the nth term of the sequence. -5, -25, -125, -625, an = -5 • 5n an = 5 • -5n + 1 an = 5 • -5n an = -5 • 5n - 1

OpenStudy (phi):

Your first sequence 60, 10, 5/8 , 5/18 , does not have a common ratio, so it is not a simple geometric sequence in other words to get from 60 to 10, multiply by 1/6 then from 10 multiply by 1/6 to get 10/6 = 5/3 but your third term is 5/8.

OpenStudy (phi):

But each term is smaller the the one previous to it. So it is not diverging (growing larger and larger). So we can rule out choices about diverging and it is definitely not approaching 72 or -15540

OpenStudy (anonymous):

@phi Converges; 0??

OpenStudy (phi):

If it continues getting smaller, (smaller and smaller fractions), it approaches 0

OpenStudy (anonymous):

@phi what about number 3 and 2

OpenStudy (phi):

For 2.) find an explicit rule for the nth term of the sequence. 1, 3, 9, 27, ... I would look at the differences: 3-1= 2 9-3= 6 27-9 = 18 that is not a constant difference. Then I would look at ratios: 3/1 = 3, 9/3 = 3; 27/9 = 3 that looks promising. To get from one term to the next, what do you multiply by ?

OpenStudy (anonymous):

an = 1 • 3n - 1 @phi

OpenStudy (phi):

You should put in ^ like this: 3^(n-1) other wise it looks like 3*n -1 which is different (and hard for me to figure out what you mean)

OpenStudy (phi):

to figure out that rule I would like the number of the term, starting at 1 1 then the value 1 1 2 1*3 (I used the rule, multiply the previous term by 3) 3 1*3*3 4 1*3*3*3 if you know about exponents, when you multiply 3*3, you can write 3^2 and 3*3*3 is 3^3 so the table looks like this 1 1 2 1*3^1 3 1*3^2 4 1*3^3 now notice the exponent on the 3 is 1 less than the number of the term 1 1*3^(1-1) 2 1*3^(2-1) 3 1*3^(3-1) 4 1*3^(4-1) n 1*3^(n-1) for term n, this is the value

OpenStudy (anonymous):

^^ @phi thank u so much what did u get for 3

OpenStudy (phi):

3. Find an explicit rule for the nth term of the sequence. -5, -25, -125, -625, what do you multiply by to get the next term?

OpenStudy (anonymous):

an = 5 • -5n + 1 @phi is it this one?

OpenStudy (phi):

when n is 1 , using that rule what do you get? \[ 5 \cdot -5^{n+1} \] replace n with 1 and you get \[ 5 \cdot -5^2 \] which means \[ 5 \cdot - (5\cdot 5) = - 5 \cdot 5 \cdot 5 = -125 \] but we want -5 as the first term (when n=1)

OpenStudy (anonymous):

a=-5 r=-25/-5=5 \[an=ar ^{n-1}\]

OpenStudy (phi):

You should try to see how to do the problem (or just memorize the formula surji posted... but that tends to go in one ear and out the other if you don't understand what is going on)